how much heat must be added to a system at 298K for the number of accessible states to increase by a factor of 10^6?
1
Expert's answer
2020-07-10T10:24:49-0400

"\frac{\Delta Q}{T}=k\ln{\frac{W_2}{W_1}}\to \Delta Q=T\cdot k\ln{\frac{W_2}{W_1}}="


"=T\cdot k\ln{\frac{W_2}{W_1}}=298\cdot1.38\cdot10^{-23}\cdot \ln{10^6}=5.68\cdot10^{-20}J"